OpenStudy (anonymous):

TRUE or FALSE? Polar equations can describe graphs as functions, even when their equations in the rectangular coordinate system are not functions.

OpenStudy (anonymous):

true

OpenStudy (anonymous):

True...

OpenStudy (anonymous):

Ohh, thank you! Any way you can explain it in the simplest terms?

OpenStudy (anonymous):

@ilfy214 I'll give you an example. Consider the polar equation \(\bf r = 6\) or \(\bf r=cos(\theta)+sin(\theta)\). Both equations describe circles that are functions of \(\bf \theta\). But if these same circles were to be represented in rectangular coordinates, they would have to be described implicitly and they won't be be described with y as a function of x.

OpenStudy (anonymous):

Gotcha! Thank you both :)